OpenStudy (anonymous):

Complete the square. x^2+12x-4=0

OpenStudy (anonymous):

x^2+12x+4=0 x=[-12+-sqrt(144-16)]/2=[-12+-sqrt(128… 8*sqrt2]/2=-6+-4sqrt2 x1=-6+4sqrt2 x2=-6-4sqrt2 Hope it helps!

OpenStudy (zzr0ck3r):

(x+6)^2 = 4+36 = 40 do you see what I did there?